OpenStudy (anonymous):

-9x +5y=8 7X - 4Y=0 SOLVE the system of equations using the inverse of the coefficient matrix

OpenStudy (anonymous):

Rewriting this in matrix form looks like: |-9 5||x|=|8| |7 -4||y|=|0| multiplying by the inverse of the coefficient matrix on both sides yields (x,y) that solves this system. There is a simple formula for the inverse of a 2x2 matrix that involves the determinant and shifting and negating some numbers. If A = |a b| |c d| A^(-1) = [1/det(A)] |d -b| |-c a|

OpenStudy (anonymous):

Also, det(A) = ad-bc = -6 So given that info, we can generate A^(-1) = (-1/6) |-4 -5| |-7 -9| and |x| = (-1/6) |-4 -5| |8| = |(32/6)| |y| = |-7 -9| |0| |(56/6)|

OpenStudy (anonymous):

let me know what doesn't make sense, or if the matrices aren't readable, the equation editor won't work for me =(